# Modular invariance groups and defect McKay-Thompson series

### v1

@@ 1 Introduction
The phenomenon of Monstrous Moonshine, conjectured by Conway and Norton in [1] and eventually proved by Borcherds in [4], connects the representation theory of the Monster group $\mathcal{M}$ with the theory of modular functions. Specifically, Conway and Norton conjectured that the McKay-Thompson series $T_{g}(q)$ associated with elements $g$ of $\mathcal{M}$ are Hauptmoduln for genus zero subgroups of $SL(2,\mathbb{R})$. Frenkel, Lepowsky, and Meurman’s construction in [3] of a vertex operator algebra (VOA) with outer automorphism group $\mathcal{M}$ established a connection between moonshine and conformal field theory. This remarkable correspondence has since led to deep insights into both number theory and the theory of VOAs.
In [5], Lin and Shao extended these ideas by introducing the “defect McKay-Thompson series.” Instead of computing the partition function of the monster CFT twined by a group element $g$, they considered the insertion of (generically noninvertible) topological defect lines (TDLs) into the partition function. Lin and Shao found an example of a noninvertible TDL in the monster CFT (the Kramers-Wannier defect), and showed that the associated defect McKay-Thompson series enjoyed a genus zero invariance group in $SL(2,\mathbb{R})$, though it was not the Hauptmodul. This work left open the question of whether all defect McKay-Thompson series share the genus-zero property of their classical counterparts or whether the genus-zero property found in [5] was a coincidence due to the modular functions being invariant under subgroups of low level.
In this work, we construct a number of new defect McKay-Thompson series, using decompositions of the stress tensor of the monster CFT into those of smaller RCFTs. These decompositions, inspired by Norton’s monstralizer pairs, the work of Höhn [7] and subsequent generalizations which are explicated and generalized in [6], allow us to (non-exhaustively) catalog TDLs of the monster theory, as well as their eigenvalues, from which we may construct the defect McKay-Thompson series. We may then use the known modular properties of the smaller RCFTs appearing in our decompositions to infer the modular properties of the defect McKay-Thompson series. Doing so, we find that the genus-zero property of the classical McKay-Thompson series does not in general seem to hold for their noninvertible counterparts.
@@ 2 TDLs from monstralizers
The term “monstralizer pair” is due to Norton [8], who used it to describe subgroups of $\mathcal{M}$ which mutually centralize one another. In [6], Bae et al. used Norton’s monstralizer pairs as a clue to construct decompositions of the stress tensor of the monster CFT in terms of smaller RCFTs whose symmetry groups are those subgroups of $\mathcal{M}$ appearing in the monstralizer pair. This definition may be thought of as the “VOA uplift” of the group theoretic definition of monstralizers. More precisely, if $T(z)$ is the stress tensor of the monster CFT, and CFTs $\mathcal{A}$ and $\mathcal{B}$ have symmetry groups which mutually centralize one another in $\mathcal{M}$, and stress tensors $t^{(\mathcal{A})}(z)$ and $t^{(\mathcal{B})}(z)$ satisfying
$T(z)=t^{(\mathcal{A})}(z)+t^{(\mathcal{B})}(z),$ (2.1)
then $\mathcal{A}$ and $\mathcal{B}$ may be said to form a “monstralizer pair at the level of CFTs,” and the Virasoro primaries of $V^{\natural}$ may be organized into sets of primaries with respect to $t^{(\mathcal{A})}$ or $t^{(\mathcal{B})}$ individually. This in turn implies that the partition function of $V^{\natural}$ may be decomposed as
$J(\tau)=\sum_{i}\chi_{i}^{(\mathcal{A})}(\tau)\chi_{i}^{(\mathcal{B})}(\tau)$ (2.2)
for some (possibly not all) characters $\chi_{i}^{(\mathcal{A})}$ of $A$ and $\chi_{i}^{(\mathcal{B})}$ of $B$. The modular $S$ and $T$ matrices of $\mathcal{A}$ and $\mathcal{B}$ are related by (2.2): precisely, we must have
$S^{(\mathcal{B})}=S^{(\mathcal{A})*}\mbox{\quad and\quad}T^{(\mathcal{B})}=T^{(\mathcal{A})*}$ (2.3)
@@
in order that (2.2) be modular-invariant. A variety of monstralizer pairs of VOAs are cataloged in [6]. In each of those examples one of the CFTs is an RCFT, in fact a finite sum of minimal models.
@@
Given a monstralizer pair, suppose that the CFT $\mathcal{A}$ contains a topological defect line $\mathcal{L}^{(\mathcal{A})}$, whose action on the Virasoro primaries of $\mathcal{A}$ is
$\mathcal{L}^{(\mathcal{A})}\phi_{i}^{(A)}=L_{i}^{(\mathcal{A})}\phi_{i}^{(\mathcal{A})}$ (2.4)
@@
(all the TDLs that we will consider in this work act diagonally on the primaries of interest). Then, using (2.1) and the map between Virasoro primaries of $V^{\natural}$ and Virasoro primaries of $\mathcal{A}$ (call it $\pi$), we may construct a TDL $\mathcal{L}$ of the monster CFT by imposing that the action of $\mathcal{L}$ on a given Virasoro primary of $V^{\natural}$ (say $\phi_{i}$) is
$\mathcal{L}\phi_{i}=L_{\pi(\phi_{i})}^{(\mathcal{A})}\phi_{i}.$ (2.5)
@@
Such a line commutes with $t^{(\mathcal{B})}$, since its action on the primaries of $t^{(\mathcal{B})}$ is diagonal, and commutes with $t^{(\mathcal{A})}$ since it acts on the primaries of $t^{(\mathcal{A})}$ identically to $\mathcal{L}^{(\mathcal{A})}$, which was by supposition a topological defect of $\mathcal{A}$. Thus, the line commutes with $T(z)$, so is topological, and forms a topological defect of $V^{\natural}$. We emphasize, though, that while $\mathcal{L}$ and $\mathcal{L}^{(\mathcal{A})}$ have the same eigenvalues, they may differ in other respects – for example, $\mathcal{L}$ is generically not a Verlinde line, whether or not $\mathcal{L}^{(\mathcal{A})}$ is. At the level of characters, this construction allows us to compute the partition function of the monster CFT twined by the line $\mathcal{L}$:
$Z^{\mathcal{L}}(\tau)=\sum_{i}L_{i}^{(\mathcal{A})}\chi_{i}^{(\mathcal{A})}(\tau)\chi_{i}^{(\mathcal{B})}(\tau)=\vec{\chi}^{(\mathcal{A})}(\tau)^{\intercal}D_{\mathcal{L}}\vec{\chi}^{(\mathcal{B})}(\tau)$ (2.6)
@@
where $\vec{\chi}^{(\mathcal{A})}(\tau)$ is a vector formed of those characters of $\mathcal{A}$ which appear in the monstralizer decomposition (2.2), and $D_{\mathcal{L}}=\operatorname{\text{diag}}(L_{i}^{(\mathcal{A})})$. The objects $Z^{\mathcal{L}}(\tau)$ are known as “defect McKay-Thompson series,” a term defined in [5]. In that paper, Lin and Shao identified two TDLs of the monster theory, one corresponding to the $\mathbb{Z}_{2\text{A}}$ conjugacy class of the monster and one non-invertible line, analogous to the Kramers-Wannier defect of the Ising CFT. Using the monstralizer pair $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$ in the notation of [6], we recover the same lines and the associated defect McKay-Thompson series (given in detail in section 2.1). We also consider the monstralizer pairs $\quantity(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$, $\quantity(\mathbb{Z}_{4\text{A}},4.2^{22}.Co_{3})$, and $\quantity(D_{3\text{A}},Fi_{23})$. In each case we discover new TDLs of the monster theory, and in all but one we are able to show that we have found all TDLs arising from the monstralizer pair under consideration.
@@
The monstralizer pair $(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$ has been known at the level of CFTs since [7], though in different language. In that work, Höhn constructed a subVOA of $V^{\natural}$ whose automorphism group is the baby monster $\mathbb{B}$. Lin and Shao provided a physical explanation in [5]: the fermionization of the monster CFT under an element of the $\mathbb{Z}_{2\text{A}}$ conjugacy class is the direct product of Höhn’s baby monster VOA with a single Majorana-Weyl fermion. Conversely, under bosonization, they showed that the $\mathbb{Z}_{2}$ symmetry of the direct product theory corresponding to fermion parity becomes the non-invertible element of a fusion subcategory of the monster which is isomorphic to the Ising category. In our language, this subcategory arises from the monstralizer pair $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$, which uplift to VOAs as the Ising CFT and the baby monster CFT. The Verlinde lines of the Ising CFT exhaust the fusion category of TDLs, and are reproduced (along with their action on the primaries) for convenience in table 1. Note that in this table, as well as those to follow, the top row does not label individual primaries of the monster; rather it should be understood to label sets of Virasoro primaries of $V^{\natural}$ with respect to the Ising stress tensor contained within the monster stress tensor (cf. eq 2.1).
$1$ $\sigma$ $\epsilon$
$\eta$ $1$ $1$ $-1$
$\mathcal{D}$ $\sqrt{2}$ $-\sqrt{2}$ 00
Table 1: TDLs of the Ising subcategory within the monster CFT. We emphasize that the labels $1$, $\sigma$, $\epsilon$, which refer to individual primaries in the constext of the Ising CFT, here refer to sets of Virasoro primaries of the moonshine module which are also Virasoro primaries under the Ising factor in the decomposition of the monster stress tensor.
@@
where $\quantity{1,\sigma,\epsilon}$ refer to the Ising primaries with $h=\quantity{0,\frac{1}{16},\frac{1}{2}}$ respectively, and the baby monster characters $\chi^{(V\mathbb{B}^{\natural})}_{i}$ are related to the Ising characters by the Hecke operator for vector-valued modular forms $\mathsf{T}_{47}$ [6, 9]. To obtain the defect McKay-Thompson series, we make use of equation 2.6; we find $q$-expansions for the series $Z^{\eta}$ and $Z^{\mathcal{D}}$ which agree with those in [5]. The line $\eta$ simply implements the $\mathbb{Z}_{2\text{A}}$ global symmetry of the monster, so $Z^{\eta}$ is a “typical” McKay-Thompson series, of the type studied by Borcherds, and satisfies the genus zero property. On the other hand, the line $\mathcal{D}$ is non-invertible, so $Z^{\mathcal{D}}$ is not of this type, and is not guaranteed to be genus zero.
2.2 $\quantity(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$: Three-state Potts model
@@
The next monstralizer pair we consider is $(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$, which uplift to VOAs as the three-state Potts model and its commutant in $V^{\natural}$. Unlike the Ising case, the commutant VOA here is not a familiar one (though it was studied in [10]); we shall only need its characters, which are images of the three-state Potts characters under the action of the Hecke operator $\mathsf{T}_{29}$. Also unlike the Ising case, the Verlinde lines of the three-state Potts model do not exhaust its fusion category: there is another non-invertible line which commutes with the Virasoro algebra, but not the extended chiral algebra [11, 12]. The TDLs in this category and their eigenvalues are given in table 2.
$(l,m)$ $(3,3)$ $(3,1)$ $(3,-1)$ $(2,0)$ $(1,1)$ $(2,2)$
$\eta$ $1$ $-\alpha$ $\alpha^{2}$ $1$ $-\alpha$ $\alpha^{2}$
$\eta^{2}$ $1$ $\alpha^{2}$ $-\alpha$ $1$ $\alpha^{2}$ $-\alpha$
$W$ $\varphi$ $\varphi$ $\varphi$ $-\varphi^{-1}$ $-\varphi^{-1}$ $-\varphi^{-1}$
$\eta W$ $\varphi$ $-\alpha\varphi$ $\alpha^{2}\varphi$ $-\varphi^{-1}$ $-\alpha\varphi^{-1}$ $-\alpha^{2}\varphi^{-1}$
$\eta^{2}W$ $\varphi$ $\alpha^{2}\varphi$ $-\alpha\varphi$ $-\varphi^{-1}$ $-\alpha^{2}\varphi^{-1}$ $-\alpha\varphi^{-1}$
$N$ $\sqrt{3}$ 00 00 $-\sqrt{3}$ 00 00
$WN$ $\varphi\sqrt{3}$ 00 00 $\varphi^{-1}\sqrt{3}$ 00 00
@@
Table 2: The fusion subcategory of the monster CFT inherited from the three-state Potts model. Here $\alpha=(-1)^{1/3}$, and $\varphi=\frac{1+\sqrt{5}}{2}$.
@@
We note that the only characters appearing in the monstralizer decomposition of $J$ in this case correspond to the chiral primaries of the three-state Potts model (rather than the Virasoro primaries), and so the TDL eigenvalues are only given for these. The lines labeled $W$, $\eta W$, $\eta^{2}W$, $N$, and $WN$ are non-invertible. Once again, using the Hecke operator for vector-valued modular forms and equation 2.6, we are able to compute $q$-expansions for the defect McKay-Thompson series associated to each line. $Z^{\mathcal{\eta}}$ and $Z^{\mathcal{\eta}^{2}}$ are the usual McKay-Thompson series for the $\mathbb{Z}_{3A}$ conjugacy class; the other series are novel.

### v2

@@ 1 Introduction
The phenomenon of monstrous moonshine, conjectured by Conway and Norton in [1], connects the representation theory of the Monster group $\mathcal{M}$ with the theory of modular functions. Specifically, Conway and Norton conjectured that the McKay-Thompson series $T_{g}(q)$ associated with elements $g$ of $\mathcal{M}$ are Hauptmoduln for genus zero subgroups of $\operatorname{\textsl{SL}}(2,\mathbb{R})$. Frenkel, Lepowsky, and Meurman’s construction in [3] of a vertex operator algebra (VOA) with outer automorphism group $\mathcal{M}$ established a connection between moonshine and conformal field theory. The genus zero property was then proved by Borcherds [4]. This remarkable correspondence has since led to deep insights into both number theory and the theory of VOAs.
In [5], Lin and Shao extended these ideas by introducing the “defect McKay-Thompson series.” Instead of computing the partition function of the monster CFT twined by a group element $g$, they considered the insertion of (generically non-invertible) topological defect lines (TDLs) into the partition function. Lin and Shao found an example of a non-invertible TDL in the monster CFT (the Kramers-Wannier defect), and showed that the associated defect McKay-Thompson series enjoyed a genus zero invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{R})$, though it was not the Hauptmodul. This work left open the question of whether all defect McKay-Thompson series share the genus-zero property of their classical counterparts or whether the genus-zero property found in [5] was a coincidence due to the invariance of the modular functions under $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ subgroups of low level.
If $\Gamma$ is a discrete subgroup of $\operatorname{\textsl{SL}}(2,\mathbb{R})$ that is commensurable with $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, then the quotient of the upper half plane $\mathbb{H}$ (plus cusps) by $\Gamma$ is a compact Riemann surface. If the surface has genus zero, then we say that $\Gamma$ is a genus zero subgroup of $\operatorname{\textsl{SL}}(2,\mathbb{R})$. The genus zero property of the McKay-Thompson series of monstrous moonshine is one of the most striking and mysterious aspects of monstrous moonshine. Genus zero groups also play a central role in organizing other instances of moonshine including umbral and penumbral moonshine [7, 8]. Thus, it is of great interest to determine whether or not there is an extension of the genus zero property to defect McKay-Thompson series.
In this work, we construct a number of new defect McKay-Thompson series, using decompositions of the stress tensor of the monster CFT into those of smaller RCFTs. These decompositions, inspired by Norton’s monstralizer pairs, the work of Höhn [9] and subsequent generalizations which are explicated and generalized in [6], allow us to (non-exhaustively) catalog TDLs of the monster theory, as well as their eigenvalues, from which we may construct the defect McKay-Thompson series. We may then use the known modular properties of the smaller VOAs appearing in our decompositions to infer the modular properties of the defect McKay-Thompson series. Doing so, we find that the genus-zero property of the classical McKay-Thompson series does not hold in general for their non-invertible counterparts.
@@ 2 TDLs from monstralizers
The term “monstralizer pair” is due to Norton [10], who used it to describe subgroups of $\mathcal{M}$ which mutually centralize one another. In [6], Bae et al. used Norton’s monstralizer pairs as a clue to construct decompositions of the stress tensor of the monster CFT in terms of commuting conformal vectors for chiral subalgebras of $V^{\natural}$ whose symmetry groups are those subgroups of $\mathcal{M}$ appearing in the monstralizer pair. This definition may be thought of as the “VOA uplift” of the group-theoretic definition of monstralizers. More precisely: say the monster module $V^{\natural}$ has subVOA (chiral subalgebra) $\mathcal{A}$ with commutant $\mathcal{B}$, the outer automorphism groups of which form a monstralizing pair of subgroups in the sense of Norton. This implies that the partition function of $V^{\natural}$ may be decomposed as
$J(\tau)=\sum_{ij}N_{ij}\chi_{i}^{(\mathcal{A})}(\tau)\chi_{j}^{(\mathcal{B})}(\tau)$ (2.1)
for some (possibly not all) characters $\chi_{i}^{(\mathcal{A})}$ of $\mathcal{A}$ and $\chi_{i}^{(\mathcal{B})}$ of $\mathcal{B}$; moreover, the conformal vectors $t_{\mathcal{A}}(z)$ and $t_{\mathcal{B}}(z)$ associated to these chiral subalgebras must combine to form the $c=24$ stress tensor of the monster CFT:
$T(z)=t_{\mathcal{A}}(z)+t_{\mathcal{B}}(z).$ (2.2)
We then say that $\mathcal{A}$ and $\mathcal{B}$ form a monstralizer pair at the level of VOAs. Next, we seek out a (modular invariant) RCFT, which we call the monstralizing RCFT, whose chiral algebra is $\mathcal{A}$. By modular invariance of Klein’s $j$ invariant, the antichiral algebra $\bar{\mathcal{A}}^{\prime}$ of the RCFT must share modular data with $\mathcal{B}$:
$S^{(\mathcal{B})}=S^{(\bar{\mathcal{A}}^{\prime})}=S^{(\mathcal{A})*}\mbox{\quad and\quad}T^{(\mathcal{B})}=T^{(\bar{\mathcal{A}}^{\prime})}=T^{(\mathcal{A})*},$ (2.3)
@@
and we demand that, upon matching $\mathcal{A}\otimes\mathcal{B}$ primaries in $V^{\natural}$ with $\mathcal{A}\otimes\bar{\mathcal{A}}^{\prime}$ primaries in the monstralizing RCFT, the multiplicities $N_{ij}$ are the same. A variety of monstralizer pairs of VOAs are cataloged in [6]. In each of those examples the monstralizing RCFT is a finite sum of minimal models.
@@
Given a monstralizer pair of VOAs, suppose that the monstralizing RCFT contains a topological defect line $\mathcal{L}^{(\mathcal{A})}$ (preserving $\mathcal{A}\otimes\bar{\mathcal{A}}^{\prime}$), whose action on the primaries is
$\mathcal{L}^{(\mathcal{A})}\phi_{ij}^{(A)}=L_{ij}^{(\mathcal{A})}\phi_{ij}^{(\mathcal{A})}$ (2.4)
@@
where $(i,j)$ are the left- and right-moving indices from (2.1). More generally, topological defects can act non-diagonally on the primaries, but this type of TDL will not appear in our analysis. Then, using the map between $\mathcal{A}\otimes\mathcal{B}$ primaries of $V^{\natural}$ and $\mathcal{A}\otimes\bar{\mathcal{A}}^{\prime}$ primaries of the monstralizing RCFT (call it $\pi$), we may construct a TDL $\mathcal{L}$ of the monster CFT by imposing that the action of $\mathcal{L}$ on a given $\mathcal{A}\otimes\mathcal{B}$ primary of $V^{\natural}$ (say $\phi_{i}$) is
$\mathcal{L}\phi_{ij}=L_{\pi(\phi_{ij})}^{(\mathcal{A})}\phi_{ij}.$ (2.5)
@@
Such a line is topological by construction (acting identically on primaries and descendants), and (2.1) and (2.3) guarantee that it satisfies the Cardy condition since $\mathcal{L}^{(\mathcal{A})}$ does so; it therefore forms a valid TDL in the monster CFT. Moreover, $\mathcal{L}$ will preserve the subalgebra $\mathcal{A}\otimes\mathcal{B}$ of $V^{\natural}$. In fact, we need not start with a line preserving the maximal left-moving chiral algebra of the monstralizing RCFT. More generally, applying (2.5) to a defect preserving $\mathcal{A}^{\prime}\otimes\bar{\mathcal{A}}^{\prime}$ in the monstralizing RCFT (with Virasoro $\mathcal{A}^{\prime}<\mathcal{A}$) yields a TDL of the monster CFT preserving $\mathcal{A}^{\prime}\otimes\mathcal{B}$. We emphasize, though, that while $\mathcal{L}$ and $\mathcal{L}^{(\mathcal{A})}$ have the same eigenvalues, they may differ in other respects – for example, $\mathcal{L}$ is generically not a Verlinde line, whether or not $\mathcal{L}^{(\mathcal{A})}$ is. At the level of characters, this construction allows us to compute the partition function of the monster CFT twined by the line $\mathcal{L}$:
$Z^{\mathcal{L}}(\tau)=\sum_{i}L_{ij}^{(\mathcal{A})}N_{ij}\chi_{i}^{(\mathcal{A})}(\tau)\chi_{j}^{(\mathcal{B})}(\tau)\equiv\vec{\chi}^{(\mathcal{A})}(\tau)^{\intercal}D_{\mathcal{L}}\vec{\chi}^{(\mathcal{B})}(\tau)$ (2.6)
@@
where $\vec{\chi}^{(\mathcal{A})}(\tau)$ (resp. $\vec{\chi}^{(\mathcal{B})}(\tau)$) is a vector formed of those characters of $\mathcal{A}$ (resp. $\mathcal{B}$) which appear in the monstralizer decomposition (2.1). The objects $Z^{\mathcal{L}}(\tau)$ are known as “defect McKay-Thompson series,” a term defined in [5]. In that paper, Lin and Shao identified two TDLs of the monster theory, one corresponding to the $\mathbb{Z}_{2\text{A}}$ conjugacy class of the monster and one non-invertible line, analogous to the Kramers-Wannier defect of the Ising CFT. Using the monstralizer pair $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$ in the notation of [6], we recover the same lines and the associated defect McKay-Thompson series (given in detail in section 2.1). We also consider the monstralizer pairs $\quantity(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$, $\quantity(\mathbb{Z}_{4\text{A}},4.2^{22}.Co_{3})$, and $\quantity(D_{3\text{A}},Fi_{23})$. In each case we discover new TDLs of the monster theory, and in two cases we are able to show that we have found all TDLs arising from the monstralizer pair under consideration.
@@
The monstralizer pair $(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$ has been known at the level of VOAs since [9], though in different language. In that work, Höhn constructed a subVOA of $V^{\natural}$ whose automorphism group is the baby monster $\mathbb{B}$. Lin and Shao provided a physical explanation in [5]: the fermionization of the monster CFT under an element of the $\mathbb{Z}_{2\text{A}}$ conjugacy class is the direct product of Höhn’s baby monster VOA with a single Majorana-Weyl fermion. Conversely, under bosonization, they showed that the $\mathbb{Z}_{2}$ symmetry of the direct product theory corresponding to fermion parity becomes the non-invertible element of a fusion subcategory of the monster which is isomorphic to the Ising category. In our language, this subcategory arises from the monstralizer pair $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$, which uplift to VOAs as Virasoro at $c=\frac{1}{2}$ and the baby monster VOA; we may take the associated monstralizing RCFT to be the Ising CFT. The Verlinde lines of the Ising CFT exhaust the fusion category of TDLs - i.e. no other Virasoro-preserving lines exist. They are reproduced (along with their action on the primaries) for convenience in table 1. Note that in this table, as well as those to follow, the top row does not label individual Virasoro primaries of the monster; rather it should be understood to label sets of primaries of $V^{\natural}$ with respect to the “Ising” chiral subalgebra.
11 $\sigma$ $\epsilon$
11 11 11 11
$\eta$ 11 11 1-1
$\mathcal{D}$ $\sqrt{2}$ $\sqrt{2}$ 0
Table 1: TDLs of the Ising subcategory within the monster CFT. We emphasize that the labels 11, $\sigma$, $\epsilon$, which refer to Virasoro primaries in the constext of the Ising CFT, here refer to primaries of $V^{\natural}$ under its $c=\frac{1}{2}$ Virasoro chiral subalgebra.
@@ 11 $\sigma$ $\epsilon$
where $\quantity{1,\sigma,\epsilon}$ refer to the Ising primaries with $h=\quantity{0,\frac{1}{16},\frac{1}{2}}$ respectively, and the baby monster characters $\chi^{(V\mathbb{B}^{\natural})}_{i}$ are related to the Ising characters by the Hecke operator for vector-valued modular forms $\mathsf{T}_{47}$ [6, 11]. To obtain the defect McKay-Thompson series, we make use of equation 2.6; we find $q$-expansions for the series $Z^{\eta}$ and $Z^{\mathcal{D}}$ which agree with those in [5]. The line $\eta$ simply implements the $\mathbb{Z}_{2\text{A}}$ global symmetry of the monster, so $Z^{\eta}$ is a “typical” McKay-Thompson series, of the type studied by Borcherds, and satisfies the genus zero property. On the other hand, the line $\mathcal{D}$ is non-invertible, so $Z^{\mathcal{D}}$ is not of this type, and is not guaranteed to possess a genus zero invariance group.
2.2 $\quantity(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$: $\mathbb{Z}_{3}$ parafermion
@@
The next monstralizer pair we consider is $(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$, which uplift to VOAs as the chiral algebra $W_{3}$ and its commutant in $V^{\natural}$, with monstralizing RCFT the $\mathbb{Z}_{3}$ parafermion CFT. Unlike the $\mathbb{Z}_{2A}$ case, the commutant VOA here is not a familiar one (though it was studied in [12]); we shall only need its characters, which are images of the parafermion characters under the action of the Hecke operator $\mathsf{T}_{29}$. Also unlike the $\mathbb{Z}_{2A}$ case, the Verlinde lines of the $\mathbb{Z}_{3}$ parafermion CFT do not exhaust its fusion category of topological defects: there are additional lines which commute with the Virasoro algebra at $c=\frac{4}{5}$, but not the extended chiral algebra [13, 14]. Per the discussion in the introduction to this section, to obtain as many monstrous TDLs as possible we should start with TDLs of the $\mathbb{Z}_{3}$ parafermion CFT preserving $(\text{Virasoro}\otimes\overline{W}_{3})$. The TDLs in this category and their eigenvalues are given in table 2.
11 $Y$ $\Omega$ $\tilde{\Omega}$ $\epsilon$ $X$ $\Phi$ $\tilde{\Phi}$ $Z$ $Z^{*}$ $\sigma$ $\sigma^{*}$ 11 11 11 11 11 11 11 11 11 11 11 11 11 $\eta$ 11 11 11 11 11 11 11 11 $\alpha$ $\alpha^{2}$ $\alpha$ $\alpha^{2}$ $\eta^{2}$ 11 11 11 11 11 11 11 11 $\alpha^{2}$ $\alpha$ $\alpha^{2}$ $\alpha$ $W$ $\varphi$ $\varphi$ $\varphi$ $\varphi$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi$ $\varphi$ $\varphi^{-1}$ $\varphi^{-1}$ $\eta W$ $\varphi$ $\varphi$ $\varphi$ $\varphi$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$ $\alpha\varphi$ $\alpha^{2}\varphi$ $\alpha\varphi^{-1}$ $\alpha^{2}\varphi^{-1}$ $\eta^{2}W$ $\varphi$ $\varphi$ $\varphi$ $\varphi$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi^{-1}$ $\alpha^{2}\varphi$ $\alpha\varphi$ $\alpha^{2}\varphi^{-1}$ $\alpha\varphi^{-1}$ NN $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ 0 0 0 0 WNWN $\sqrt{3}\varphi$ $\sqrt{3}\varphi$ $\sqrt{3}\varphi$ $\sqrt{3}\varphi$ $\sqrt{3}\varphi^{-1}$ $\sqrt{3}\varphi^{-1}$ $\sqrt{3}\varphi^{-1}$ $\sqrt{3}\varphi^{-1}$ 0 0 0 0
@@
Table 2: The fusion subcategory of the monster CFT inherited from the $\mathbb{Z}_{3}$ parafermion. Here $\alpha=(-1)^{2/3}$, and $\varphi=\frac{1+\sqrt{5}}{2}$. The primaries are labeled by the corresponding Virasoro primaries of the $\mathbb{Z}_{3}$ parafermion CFT, in agreement with the labels in [14, 13].
The lines labeled $W$, $\eta W$, $\eta^{2}W$, $N$, and $WN$ are non-invertible. We pause to note that, since the modular invariant of the theory is not diagonal (with respect to Virasoro), the matrices $D_{\mathcal{L}}$ of (2.6) representing the partition function of the theory twined by the various TDLs will also not be diagonal. In particular, given a $\mathbb{Z}_{3}$ parafermion TDL $\mathcal{L}$ whose eigenvalues on the Virasoro primaries are denoted $L^{\mathcal{L}}_{\phi}$, the partition function for the monster CFT twined by the associated TDL is, using (2.6),
$Z^{\mathcal{L}}(\tau)=\vec{\chi}^{(W_{3})}(\tau)^{\intercal}\smallmatrixquantity(L^{\mathcal{L}}_{I}&L^{\mathcal{L}}_{\Omega}&&&&&&&&\\ L^{\mathcal{L}}_{\tilde{\Omega}}&L^{\mathcal{L}}_{Y}&&&&&&&&\\ &&L^{\mathcal{L}}_{\epsilon}&L^{\mathcal{L}}_{\Phi}&&&&&&\\ &&L^{\mathcal{L}}_{\tilde{\Phi}}&L^{\mathcal{L}}_{X}&&&&&&\\ &&&&L^{\mathcal{L}}_{Z}+L^{\mathcal{L}}_{Z^{*}}&&&&&\\ &&&&&L^{\mathcal{L}}_{\sigma}+L^{\mathcal{L}}_{\sigma^{*}}&&&&\\ &&&&&&0&&&\\ &&&&&&&0&&\\ &&&&&&&&0&\\ &&&&&&&&&0)\vec{\chi}^{(\tilde{W}_{3})}(\tau)$ (2.8)
@@
where $\vec{\chi}^{(W_{3})}(\tau)$ are the $\mathbb{Z}_{3}$ parafermion characters and $\vec{\chi}^{(\tilde{W}_{3})}$ are those of the commutant VOA (cf. equation 4.20 of [13]). Once again, using the Hecke operator for vector-valued modular forms, we are able to compute $q$-expansions for the defect McKay-Thompson series associated to each line. $Z^{\mathcal{\eta}}$ and $Z^{\mathcal{\eta}^{2}}$ are the usual McKay-Thompson series for the $\mathbb{Z}_{3A}$ conjugacy class; the other series are novel.